OpenStudy (anonymous):

Find all solutions of the equation on the intervals [0,2pi). 2sin^2 x+2 sin x-3=0

OpenStudy (anonymous):

what math are you in?

OpenStudy (anonymous):

per-calculs

OpenStudy (anonymous):

oh i'm only in Algebra 1 sorry :/ maybe @Notamathgenius can help you tho :)

OpenStudy (anonymous):

(don't let the name trick you ) lol

OpenStudy (anonymous):

@ciecey ?? @Notamathgenius ?

OpenStudy (notamathgenius):

substituting z for sin(x), you get: z^2 - 2z - 3 = 0 (z-3)(z+1)=0 so z, or sin(x), must either equal 3 or -1 the range of sin(x) is -1 to 1 so 3 is out of the question. Thus, setting sin(x) = -1, you get 3pi/2 radians, or 270 degree

OpenStudy (anonymous):

thank u @Notamathgenius

OpenStudy (notamathgenius):

You're welcome

OpenStudy (anonymous):

good job :D)